OpenStudy (anonymous):

Which is the incorrect number in the following series? 5, 10, 29, 84, 247

Parth (parthkohli):

Hello, I hate to say but there is no incorrect sequence/series. A sequence/series is just a bunch of numbers. Your question is wrong unless you provide what kind of sequence/series this is.

OpenStudy (anonymous):

No.. the catch is, there is a series that is to be figured out... and a nunber that is the odd one in the series.. Im juz not able to find out wch one is the odd one...

Parth (parthkohli):

Oh, got it.

Parth (parthkohli):

Look:\[10 \times 3 - 1 = 29 \]\[29\times 3 - 3 = 84\]\[84 \times 3 - 5 =247\]

Parth (parthkohli):

Look:\[10 \times 3 - 1 = 29 \]\[29\times 3 - 3 = 84\]\[84 \times 3 - 5 =247\]

Parth (parthkohli):

Looking at that pattern, the first term shold be \(3\).

Parth (parthkohli):

Looking at that pattern, the first term shold be \(3\).

Parth (parthkohli):

\[3 \times 3 - (-1) = 10\]

OpenStudy (hba):

Well i used a diff method to figure this out and it seems like 5 is the incorrect digit.

Parth (parthkohli):

What was your method? And good to hear that I am not wrong :P

OpenStudy (hba):

I am afraid my observation may seem not to be correct but i still would like to share it, 29/10=2.9 84/29=2.9 247/84=2.9 but, 10/5=2 Therfore,5 is the incorrect digit I am sorry if i was wrong .

Parth (parthkohli):

Arrey haan! Nice observation!

OpenStudy (hba):

Haha,thanks.

OpenStudy (anonymous):

Thanks to both of you... i got 2 different solution...

OpenStudy (hba):

You're welcome. But,@ParthKohli is better.

Parth (parthkohli):

Actually @hba no!

OpenStudy (hba):

I know i am wrong.

Parth (parthkohli):

All the numbers except \(\dfrac{29}{10}\) are not \(2.9\) exactly. :\

Parth (parthkohli):

Anyway, I gotta go. Good day all of you =)

OpenStudy (hba):

That is why i was calling it a wrong method. Bye.